# Closed-Form Training Dynamics Reveal Learned Features and Linear Structure in Word2Vec-like Models

### v2

@@ 1 Introduction
Figure 1: Quadratic word embedding models are a faithful and analytically solvable proxy for word2vec. We compare the time course of learning in QWEMs (top) and word2vec (bottom), finding striking similarities in their training dynamics and learned representations. Analytically, we solve for the optimization dynamics of QWEMs under gradient flow from small initialization, revealing discrete, rank-incrementing learning steps corresponding to stepwise decreases in the loss (top left). In latent space (right side plots), embedding vectors expand into subspaces of increasing dimension at each learning step. These PCA directions are the model’s learned features, and they can be extracted from our theory in closed form given only the corpus statistics and hyperparameters (1). Empirically, QWEMs yield high-quality embeddings very similar to word2vec’s in terms of their learned features and performance on benchmarks (Figure 3). See Appendix A for details.
@@
In this work, we give a closed-form description of representation learning in models trained to minimize the quartic Maclaurin approximation of the word2vec loss. We prove that the learning problem reduces to matrix factorization with quadratic loss (1), so we call these models quadratic word embedding models (QWEMs). We derive analytic solutions for their training dynamics and final embeddings under gradient flow and vanishing initialization (Figure 2). Training on Wikipedia, we show that QWEMs closely match word2vec-trained models in their training dynamics, learned features, and performance on standard benchmarks (Figure 3). We apply our results to show that the dynamical formation of abstract linear representations is well-described by quantities taken from random matrix theory (Figure 4). Taken together, our results give a clear picture of the learning dynamics in contrastive word embedding algorithms such as word2vec.
@@ 2 Preliminaries
In word2vec “skip-gram with negative sampling” (SGNS), [^1] Throughout this paper, we use the abbreviated “word2vec” to refer to the word2vec SGNS algorithm. two embedding matrices ${\bm{W}}$ and ${\bm{W}}^{\prime}$ are trained222 ${\bm{W}}$ is for “center” words, and ${\bm{W}}^{\prime}$ is for “context” words. For simplicity, we consider the setting ${\bm{W}}^{\prime}={\bm{W}}$; in Figure 3 we show that this is sufficient for good performance on semantic understanding tasks. using stochastic gradient descent to minimize the contrastive loss
@@
The averages are estimated by drawing samples from the corpus. The nonnegative hyperparameters $\{\Psi^{+}_{ij}\}$ and $\{\Psi^{-}_{ij}\}$ are reweighting coefficients for word pairs; we use them here to capture the effect of several of word2vec’s implementation details, including subsampling (i.e., probabilistically discarding frequent words during iteration), dynamic window sizes, and different negative sampling distributions. All of these can be seen as preprocessing techniques that directly modify the unigram and skip-gram distributions. In Section A.3 we provide more detail about these engineering tricks and discuss how one can encode their effects in $\Psi^{+}$ and $\Psi^{-}$.
@@ 3 Quadratic Word Embedding Models
We start by introducing a matrix ${\bm{M}}^{\star}$. We will show in 1 that ${\bm{M}}^{\star}$ is the optimization target for QWEM, just as the PMI matrix is the optimization target for word2vec.
@@
To understand this quantity, first note that if language were a stochastic process with independently sampled words, the co-occurrence statistics would be structureless, i.e., $P_{ij}-P_{i}P_{j}=0$. The distributional hypothesis then suggests that algorithms may learn semantics by modeling the statistical deviations from independence. It is exactly these (relative) deviations that comprise the optimization target ${\bm{M}}^{\star}$ and serve as the central statistics of interest in our theory. Furthermore, ${\bm{M}}^{\star}$ can be seen as an approximation the PMI matrix; see Appendix D.
@@
Our goal now is to directly convert Equation 3 into a matrix factorization problem. To do this, we make some judicious choices for the hyperparameters. We first define the quantity ${\bm{G}}_{ij}\coloneqq\Psi^{+}_{ij}P_{ij}+\Psi^{-}_{ij}P_{i}P_{j}$, which captures the aggregate effect of the reweighting hyperparameters on the optimization. Then we establish the following hyperparameter setting.
@@
Note that infinitely many choices of $\Psi^{+}$ and $\Psi^{-}$ are encompassed by this setting. Let us study a concrete example: $\Psi^{+}_{ij}=\Psi^{-}_{ij}=(P_{ij}+P_{i}P_{j})^{-1}$, so that ${\bm{G}}_{ij}=g=1$. This has the effect of down-weighting frequently appearing words and word pairs, which hastens optimization and prevents the model from over-allocating fitting power to words such as “the” or “and” which may not individually carry much semantic content. This is exactly the justification given for subsampling in word2vec, which motivates 1. Indeed, in Section A.3 we discuss how these choices of $\Psi^{+}$ and $\Psi^{-}$ can be seen as approximating several of the implementation details and engineering tricks in word2vec. In Figures 2 and 3, we show that this simplified hyperparameter setting does not wash out the relevant structure, thus retaining realism.
With these definitions, we state our key result: rank-constrained quadratic word embedding models trained under 1 learn the top $d$ eigendirections of ${\bm{M}}^{\star}$.
@@
[QWEM = unweighted matrix factorization]theoremmatrixfac Under 1, the contrastive loss Equation 3 can be rewritten as the unweighted matrix factorization problem
@@
Proof. Equation 6 follows from completing the square in Equation 3. Equation 7 follows from the Eckart-Young-Mirsky theorem. We note that the assumption that ${\bm{\Lambda}}_{[:d,:d]}$ is PSD is usually easily satisfied in practice (see Section A.2).
@@
The main limitation of 1 is its restriction to 1. What happens in the general case?
@@
propositionweightedmfac Equation 3 can be rewritten as the weighted matrix factorization problem
@@
Proof. See Appendix C. Equation 8 says that in the general case, the form of the target matrix remains unchanged. Equation 9 says that if the hyperparameters are chosen so that ${\bm{G}}$ is rank-1, then we can characterize the minimizer. However, if ${\bm{G}}$ is not rank-1, then we do not know what low-rank factorization is learned, nor can we describe the training dynamics; indeed, weighted matrix factorization is known to be NP-hard (Gillis & Glineur, 2011). For this reason, we do not revisit these more general settings, and we hereafter restrict our focus to 1. This is ultimately justified by the empirically close match between our theory and the embeddings learned by word2vec (Figure 3).
@@
Note that despite the simplification afforded by 1, the minimization problem Equation 6 is still nonconvex since $\mathcal{L}({\bm{W}})$ is quartic in ${\bm{W}}$. Thus, there remain questions regarding convergence time and the effect of early stopping regularization. To study them, we examine the training trajectories induced by gradient flow. The central variables of our theory will be the economy-sized SVD of the embeddings, ${{\bm{W}}(t)}^{\top}={\bm{U}}(t){\bm{S}}(t){{\bm{V}}(t)}^{\top}$, and the eigendecomposition of the target, ${\bm{M}}^{\star}={\bm{V}}^{\star}{\bm{\Lambda}}{{\bm{V}}^{\star}}^{\top}$. For convenience, we define $\lambda_{k}\coloneqq{\bm{\Lambda}}_{kk}$ and $s_{k}\coloneqq{\bm{S}}_{kk}$.
In general, it is difficult to solve the gradient descent dynamics of Equation 3 from arbitrary initialization. We first consider a simple toy case in which the initial embeddings are already aligned with the top $d$ eigenvectors of ${\bm{M}}^{\star}$.
@@
The proof is given in Appendix C. See Saxe et al. (2014) for a proof in an equivalent learning problem.
@@
Figure 2: Theory matches experiment. We make two simplifications to the word2vec algorithm: a quartic approximation of the loss, and a restriction on the reweighting hyperparameters. We train these QWEMs on 2 billion tokens of English Wikipedia (see Appendix A for details) and compare to word2vec. We find good qualitative match in the singular value dynamics, both with the standard word2vec initialization scheme and with small random initialization. (For evidence that the singular vectors match as well, see Figure 3.) We compare the dynamics to the prediction of Figure 2, which is derived in the vanishing initialization limit with full-batch gradient flow. Even though the experiment uses stochastic mini-batching, non-vanishing learning rate, and large initialization, we find excellent agreement even up to constant factors.
@@
[Training dynamics, vanishing random initialization]resultsri Initialize ${\bm{W}}_{ij}(0)\sim\mathcal{N}(0,\sigma^{2})$, and let ${\bm{S}}(0)$ denote the singular values of ${\bm{W}}(0)$. Define the characteristic time $\tau_{1}\coloneqq\lambda_{1}^{-1}\ln(\lambda_{1}/\sigma^{2})$ and the rescaled time variable $\tilde{t}=t/\tau_{1}$. Define ${\bm{W}}^{\star}(0)\coloneqq{\bm{V}}_{[:,:d]}^{\star}{\bm{S}}(0)$ and let ${\bm{W}}^{\star}(t)$ be its gradient flow trajectory given by Section 3.2. If ${\bm{\Lambda}}_{[:d,:d]}$ is positive semidefinite, then under 1 we have with high probability that
@@
Derivation. See Appendix C. The main idea is to study the dynamics of the ${\bm{Q}}{\bm{R}}$ factorization of ${{\bm{W}}}^{\top}{{\bm{V}}^{\star}}^{\top}$. We write the equation of motion, discard terms that become small in the limit, solve the resulting equation, and show that the discarded terms remain small. We conjecture that our argument can be made rigorous by appropriately bounding the discarded terms. We leave this to future work.
Figure 2 generalizes previous work that establishes a silent alignment phenomenon for linear networks with scalar outputs (Atanasov et al., 2022). Here, for all $k\leq d$, ${\bm{V}}_{[:,:k]}$ quickly aligns with ${\bm{V}}^{\star}_{[:,:k]}$ while $s_{k}$ remains near initialization; thus the alignment assumption is quickly near-satisfied and Section 3.2 becomes applicable. We conclude that quadratic word embedding models trained from small random initialization are inherently greedy spectral methods. The principal components of the embeddings learn a one-to-one correspondence with the eigenvectors of ${\bm{M}}^{\star}$, and each component is realized independently and sequentially. Thus, early stopping acts as an implicit regularizer, constraining model capacity in terms of rank rather than weight norm.
@@
word2vec QWEM svd(PPMI)
@@
Figure 3: QWEMs and word2vec learn similar features, whereas PMI (and variants) differ qualitatively. For the following, all models ${\bm{W}}\in\mathbb{R}^{V\times d}$ are trained with $V=10,000$ and $d=200$ on 2 billion tokens of Wikipedia. See Appendix A for experimental details. We denote $\mathrm{svd}({\bm{M}})\coloneqq\arg\min_{\bm{W}}\|{\bm{W}}{{\bm{W}}}^{\top}-{\bm{M}}\|_{\mathrm{F}}^{2}$. (Top.) We compute the principal components of the final embeddings and list the closest embeddings. See Appendix B for a quantitative plot of subspace overlaps. Top section: top three components of word2vec and QWEM represent topic-level concepts (biography, government, geography) corresponding to common topics on Wikipedia. Middle section: components closest to word2vec components 11, 12, 13. Up to reordering, QWEMs match closely and remain interpretable, while positive PMI deviates. Bottom row: late components lose their interpretability. (Bottom.) Since QWEMs and word2vec learn similar features, they perform similarly on the vector addition analogy completion task. Explicit factorization of ${\bm{M}}^{\star}$ almost matches the performance of word2vec, much better than the best previously-known SVD embeddings.
Our rationale for studying QWEMs is to gain analytical tractability (e.g., Figure 2) in a setting that is “close to” the true setting of interest: word2vec. In Figure 3, we check whether QWEMs are in fact a faithful representative for word2vec. We find that not only do QWEMs nearly match word2vec on the analogy completion task, they also learn very similar representations, as measured by the alignment between their singular vectors. Importantly, QWEMs are closer to word2vec than the least-squares approximations of either PMI or PPMI. This underscores the importance of accounting for the implicit bias of gradient descent.
To run the experiments in Figures 1, 2 and 3, we implemented a GPU-enabled training algorithm for both word2vec and QWEM. The corpus data is streamed from the hard drive in chunks to avoid memory overhead and excessive disk I/O, and the loops for batching positive and negative word pairs are compiled for efficiency. With these optimizations, both word2vec and QWEM in Figure 3 train in 75 minutes on a home desktop computer equipped with a GeForce GTX 1660 GPU. Notably, if $V$ is not too large, it is faster to construct and explicitly factorize ${\bm{M}}^{\star}$: $\sim$10 minutes in total for $V=10,000$.
Figure 3 suggests that if one is interested in understanding the learning behaviors of word2vec, it suffices to study QWEMs. Then 1 and Figure 2 state that if one is interested in understanding the learning behaviors of QWEMs, it suffices to study ${\bm{M}}^{\star}$. In this spirit, we will now study ${\bm{M}}^{\star}$ to investigate certain aspects of representation learning in word embedding models.

### v3

@@ Abstract
Code to reproduce all experiments available at https://github.com/dkarkada/qwem.
@@ 1 Introduction
Figure 1: Quadratic word embedding models are a faithful and analytically solvable proxy for word2vec. We compare the time course of learning in QWEMs (top) and word2vec (bottom), finding striking similarities in their training dynamics and learned representations. Analytically, we solve for the optimization dynamics of QWEMs under gradient flow from small initialization, revealing discrete, rank-incrementing learning steps corresponding to stepwise decreases in the loss (top left). In latent space (right side plots), embedding vectors expand into subspaces of increasing dimension at each learning step. These PCA directions are the model’s learned features, and they can be extracted from our theory in closed form given only the corpus statistics and hyperparameters (˜1). Empirically, QWEMs yield high-quality embeddings very similar to word2vec’s in terms of their learned features and performance on benchmarks (Figure˜3). See Appendix˜A for details.
@@
In this work, we give a closed-form description of representation learning in models trained to minimize the quartic Maclaurin approximation of the word2vec loss. We prove that the learning problem reduces to matrix factorization with quadratic loss (˜1), so we call these models quadratic word embedding models (QWEMs). We derive analytic solutions for their training dynamics and final embeddings under gradient flow and vanishing initialization (Figure˜2). Training on Wikipedia, we show that QWEMs closely match word2vec-trained models in their training dynamics, learned features, and performance on standard benchmarks (Figure˜3). We apply our results to show that the dynamical formation of abstract linear representations is well-described by quantities taken from random matrix theory (Figure˜4). Taken together, our results give a clear picture of the learning dynamics in contrastive word embedding algorithms such as word2vec.
@@ 2 Preliminaries
[…]
@@
The averages are estimated by drawing samples from the corpus. The nonnegative hyperparameters $\{\Psi^{+}_{ij}\}$ and $\{\Psi^{-}_{ij}\}$ are reweighting coefficients for word pairs; we use them here to capture the effect of several of word2vec’s implementation details, including subsampling (i.e., probabilistically discarding frequent words during iteration), dynamic window sizes, and different negative sampling distributions. All of these can be seen as preprocessing techniques that directly modify the unigram and skip-gram distributions. In Section˜A.3 we provide more detail about these engineering tricks and discuss how one can encode their effects in $\Psi^{+}$ and $\Psi^{-}$.
@@ 3 Quadratic Word Embedding Models
We start by introducing a matrix $\bm{M}$$^{\star}$. We will show in ˜1 that $\bm{M}$$^{\star}$ is the optimization target for QWEM, just as the PMI matrix is the optimization target for word2vec.
@@
To understand this quantity, first note that if language were a stochastic process with independently sampled words, the co-occurrence statistics would be structureless, i.e., $P_{ij}-P_{i}P_{j}=0$. The distributional hypothesis then suggests that algorithms may learn semantics by modeling the statistical deviations from independence. It is exactly these (relative) deviations that comprise the optimization target $\bm{M}$$^{\star}$ and serve as the central statistics of interest in our theory. Furthermore, $\bm{M}$$^{\star}$ can be seen as an approximation the PMI matrix; see Appendix˜D.
@@
Our goal now is to directly convert Equation˜3 into a matrix factorization problem. To do this, we make some judicious choices for the hyperparameters. We first define the quantity ${\bm{G}}_{ij}\coloneqq\Psi^{+}_{ij}P_{ij}+\Psi^{-}_{ij}P_{i}P_{j}$, which captures the aggregate effect of the reweighting hyperparameters on the optimization. Then we establish the following hyperparameter setting.
@@
Note that infinitely many choices of $\Psi^{+}$ and $\Psi^{-}$ are encompassed by this setting. Let us study a concrete example: $\Psi^{+}_{ij}=\Psi^{-}_{ij}=(P_{ij}+P_{i}P_{j})^{-1}$, so that ${\bm{G}}_{ij}=g=1$. This has the effect of down-weighting frequently appearing words and word pairs, which hastens optimization and prevents the model from over-allocating fitting power to words such as “the” or “and” which may not individually carry much semantic content. This is exactly the justification given for subsampling in word2vec, which motivates ˜1. Indeed, in Section˜A.3 we discuss how these choices of $\Psi^{+}$ and $\Psi^{-}$ can be seen as approximating several of the implementation details and engineering tricks in word2vec. In Figures˜2 and 3, we show that this simplified hyperparameter setting does not wash out the relevant structure, thus retaining realism.
With these definitions, we state our key result: rank-constrained quadratic word embedding models trained under ˜1 learn the top $d$ eigendirections of $\bm{M}$$^{\star}$.
@@
[QWEM = unweighted matrix factorization]theoremmatrixfac Under ˜1, the contrastive loss Equation˜3 can be rewritten as the unweighted matrix factorization problem
@@
Proof. Equation˜6 follows from completing the square in Equation˜3. Equation˜7 follows from the Eckart-Young-Mirsky theorem. In Section˜A.2, we note that the PSD assumption is easily satisfied in practice.
@@
The main limitation of ˜1 is its restriction to ˜1. What happens in the general case?
@@
propositionweightedmfac Equation˜3 can be rewritten as the weighted matrix factorization problem
@@
Proof. See Appendix˜C. Equation˜8 says that in the general case, the form of the target matrix remains unchanged. Equation˜9 says that if the hyperparameters are chosen so that $\bm{G}$} is rank-1, then we can characterize the minimizer. However, if $\bm{G}$} is not rank-1, then we do not know what low-rank factorization is learned, nor can we describe the training dynamics; indeed, weighted matrix factorization is known to be NP-hard (Gillis & Glineur, 2011). For this reason, we do not revisit these more general settings, and we hereafter restrict our focus to ˜1. This is ultimately justified by the empirically close match between our theory and the embeddings learned by word2vec (Figure˜3).
@@
Note that despite the simplification afforded by ˜1, the minimization problem Equation˜6 is still nonconvex since $\mathcal{L}({\bm{W}})$ is quartic in $\bm{W}$}. Thus, there remain questions regarding convergence time and the effect of early stopping regularization. To study them, we examine the training trajectories induced by gradient flow. The central variables of our theory will be the economy-sized SVD of the embeddings, ${{\bm{W}}(t)}^{\top}={\bm{U}}(t){\bm{S}}(t){{\bm{V}}(t)}^{\top}$, and the eigendecomposition of the target, ${\bm{M}}^{\star}={\bm{V}}^{\star}{\bm{\Lambda}}{{\bm{V}}^{\star}}^{\top}$. For convenience, we define $\lambda_{k}\coloneqq{\bm{\Lambda}}_{kk}$ and $s_{k}\coloneqq{\bm{S}}_{kk}$.
In general, it is difficult to solve the gradient descent dynamics of Equation˜3 from arbitrary initialization. We first consider a simple toy case in which the initial embeddings are already aligned with the top $d$ eigenvectors of $\bm{M}$$^{\star}$.
@@
The proof is given in Appendix˜C. See Saxe et al. (2014) for a proof in an equivalent learning problem.
@@
Figure 2: Theory matches experiment. We make two simplifications to the word2vec algorithm: a quartic approximation of the loss, and a restriction on the reweighting hyperparameters. We train these QWEMs on 2 billion tokens of English Wikipedia (see Appendix˜A for details) and compare to word2vec. We find good qualitative match in the singular value dynamics, both with the standard word2vec initialization scheme and with small random initialization. (For evidence that the singular vectors match as well, see Figure˜3.) We compare the dynamics to the prediction of Figure˜2, which is derived in the vanishing initialization limit with full-batch gradient flow. Even though the experiment uses stochastic mini-batching, non-vanishing learning rate, and large initialization, we find excellent agreement even up to constant factors.
@@
[Training dynamics, vanishing random initialization]resultsri Initialize ${\bm{W}}_{ij}(0)\sim\mathcal{N}(0,\sigma^{2})$, and let ${\bm{S}}(0)$ denote the singular values of ${\bm{W}}(0)$. Define the characteristic time $\tau_{1}\coloneqq\lambda_{1}^{-1}\ln(\lambda_{1}/\sigma^{2})$ and the rescaled time variable $\tilde{t}=t/\tau_{1}$. Define ${\bm{W}}^{\star}(0)\coloneqq{\bm{V}}_{[:,:d]}^{\star}{\bm{S}}(0)$ and let ${\bm{W}}^{\star}(t)$ be its gradient flow trajectory given by Section˜3.2. If ${\bm{\Lambda}}_{[:d,:d]}$ is positive semidefinite, then under ˜1 we have with high probability that
@@
Derivation. See Appendix˜C. The main idea is to study the dynamics of the ${\bm{Q}}{\bm{R}}$ factorization of ${{\bm{W}}}^{\top}{\bm{V}}^{\star}$. We write the equation of motion, discard terms that become small in the limit, solve the resulting equation, and show that the discarded terms remain small. We conjecture that our argument can be made rigorous by appropriately bounding the discarded terms. We leave this to future work.
Figure˜2 generalizes previous work that establishes a silent alignment phenomenon for linear networks with scalar outputs (Atanasov et al., 2022). Here, for all $k\leq d$, ${\bm{V}}_{[:,:k]}$ quickly aligns with ${\bm{V}}^{\star}_{[:,:k]}$ while $s_{k}$ remains near initialization; thus the alignment assumption is quickly near-satisfied and Section˜3.2 becomes applicable. We conclude that quadratic word embedding models trained from small random initialization are inherently greedy spectral methods. The principal components of the embeddings learn a one-to-one correspondence with the eigenvectors of $\bm{M}$$^{\star}$, and each component is realized independently and sequentially. Thus, early stopping acts as an implicit regularizer, constraining model capacity in terms of rank rather than weight norm.
@@
word2vec feature neighbors QWEM feature neighbors svd(PPMI) feature neighbors
@@
Figure 3: QWEMs and word2vec learn similar features, whereas PMI (and variants) differ qualitatively. For the following, all models ${\bm{W}}\in\mathbb{R}^{V\times d}$ are trained with $V=10,000$ and $d=200$ on 2 billion tokens of Wikipedia. See Appendix˜A for experimental details. We denote $\mathrm{svd}({\bm{M}})\coloneqq\arg\min_{\bm{W}}\|{\bm{W}}{{\bm{W}}}^{\top}-{\bm{M}}\|_{\mathrm{F}}^{2}$. (Top.) We compute the principal components of the final embeddings and list the closest embeddings. See Appendix˜B for a quantitative plot of subspace overlaps. Top section: top three components of word2vec and QWEM represent topic-level concepts (biography, government, geography) corresponding to common topics on Wikipedia. Middle section: components closest to word2vec components 11, 12, 13. Up to reordering, QWEMs match closely and remain interpretable, while positive PMI deviates. Bottom row: late components lose their interpretability. (Bottom.) Since QWEMs and word2vec learn similar features, they perform similarly on the vector addition analogy completion task. Explicit factorization of $\bm{M}$$^{\star}$ almost matches the performance of word2vec, much better than the best previously-known SVD embeddings.
Our rationale for studying QWEMs is to gain analytical tractability (e.g., Figure˜2) in a setting that is “close to” the true setting of interest: word2vec. In Figure˜3, we check whether QWEMs are in fact a faithful representative for word2vec. We find that not only do QWEMs nearly match word2vec on the analogy completion task, they also learn very similar representations, as measured by the alignment between their singular vectors. Importantly, QWEMs are closer to word2vec than the least-squares approximations of either PMI or PPMI. This underscores the importance of accounting for the implicit bias of gradient descent.
To run the experiments in Figures˜1, 2 and 3, we implemented a GPU-enabled training algorithm for both word2vec and QWEM. The corpus data is streamed from the hard drive in chunks to avoid memory overhead and excessive disk I/O, and the loops for batching positive and negative word pairs are compiled for efficiency. Notably, it is often faster to construct and explicitly factorize $\bm{M}$$^{\star}$ (e.g., with vocabulary size $V=10,000$, it takes $\sim$10 minutes in total on a single GeForce GTX 1660 GPU).
Figure˜3 suggests that if one is interested in understanding the learning behaviors of word2vec, it suffices to study QWEMs. Then ˜1 and Figure˜2 state that if one is interested in understanding the learning behaviors of QWEMs, it suffices to study $\bm{M}$$^{\star}$. In this spirit, we will now study $\bm{M}$$^{\star}$ to investigate certain aspects of representation learning in word embedding models.